OpenStudy (anonymous):

solve for a. 48 ≥ 6 - 7a

OpenStudy (anonymous):

\[\LARGE 48\geq 6-7a\] ad to both sides -6 \[\LARGE 48-6\geq \underbrace{6-6}_{0}-7a\] @shivam_bhalla did it already :)

OpenStudy (anonymous):

48-6 ≥6-7a-6 42 ≥-7a Multiplying by -1 on both sides, the inequality changes, −42 ≤ 7a -42/7 ≤ a a >= -6

OpenStudy (anonymous):

LOL I started to think: "should I erase my post too ?" O_O haha

OpenStudy (anonymous):

@Kreshnik , I made mistake earlier, now corrected it. Medal goes to you

OpenStudy (anonymous):

that was only a small minus. you could have fixed it easy... don't have to delete it :(

OpenStudy (anonymous):

@Kreshnik , How to edit posts?. I am 2 days old here :P

OpenStudy (anonymous):

Easy... Example: 5+3=9

OpenStudy (anonymous):

sorry 5+3=8 I made a mistake. That's it :)

OpenStudy (anonymous):

OK great. Next time I'll take care :) Thanks Kreshnik

OpenStudy (anonymous):

... Have a nice day ;).